1. 2(2x)+2x=120 solve for x
2. 4x*x=576 solve for x
3. -5(x-8)=x+28 solve for x

Answer by Cromlix(4381)

Hi there,
1) 2(2x) + 2x = 120
4x + 2x = 120
x = 120/6
x = 20
........
2) 4x*x = 576
4x^2 = 576
x^2 = 144
x = square root of 144
x = +/-12
.........
3) -5(x - 8) = x + 28
-5x + 40 = x + 28
Collect like terms
-5x - x = 28 - 40
-6x = -12
x = -12/-6
x = 2
........
Hope this helps :-)